DogGirl27:

Solve x2 – 8x = 3 by completing the square. Which is the solution set of the equation? (4 minus StartRoot 19 EndRoot comma 4 + StartRoot 19 EndRoot) (4 minus StartRoot 11 EndRoot comma 4 + StartRoot 11 EndRoot) (4 minus StartRoot 8 EndRoot comma 4 + StartRoot 8 EndRoot) (4 minus StartRoot 3 EndRoot comma 4 + StartRoot 3 EndRoot)

bivibes:

why are you asking hard questions

Alex1513:

there little hard

DogGirl27:

sorryyyy

DogGirl27:

thankss

Alex1513:

x² - 8x = 3 To complete the square add ( half the coefficient of the x- term )² to both sides x² + 2(- 4)x + 16 = 3 + 16 (x - 4)² = 19 ( take the square root of both sides ) x - 4 = ± ( add 4 to both sides ) x = 4 ± , that is x = 4 - , 4 +

Alex1513:

x = 4 ±

1 attachment